OpenStudy (anonymous):

So far this year Nick has scored a 75, 81, 88 and 92 on four of his math tests. Each test is worth a maximum of 100 points. To earn a B this year, Nick needs a test average of 80 or better. Write the inequality that represents this scenario. Let t equal Nick's grade on the fifth and the last test. _______+t≥_________

OpenStudy (tkhunny):

Have you considered calculating the average? \(\dfrac{75+81+88+92+t}{5}\ge 80\) Now, how do we find t?

OpenStudy (anonymous):

I really don't know

OpenStudy (tkhunny):

Think...think...think... Look at it and figure it out. You CAN do it.

OpenStudy (anonymous):

my school is making us do questions that my teacher has not taught us us=class

OpenStudy (tkhunny):

Do you know how to calculate an average? What is the average of 3, 5, 7? Demonstrate.

OpenStudy (anonymous):

I dontknow

OpenStudy (anonymous):

@rebeccaxhawaii

rebeccaxhawaii (rebeccaxhawaii):

well add up the numbers and divide them by 3

OpenStudy (tkhunny):

I do not believe you do not know that. Stop panicking and throwing your hands in the air. Stop and think.

rebeccaxhawaii (rebeccaxhawaii):

wow

OpenStudy (anonymous):

112

OpenStudy (anonymous):

@pooja195

pooja195 (pooja195):

@tkhunny is already helping. Listen to what he has to say he is doing a great job! :)

OpenStudy (tkhunny):

What is 112? According to Rebecca's correct response, we have \(\dfrac{3+5+7}{3} = \dfrac{15}{3} = 5 = The\;Average\) of those three numbers. It would help a lot if you ever have seen this sort of thing. Admittedly, if you are taking an online class, you may have missed it. I have not ever seen a live teacher miss the idea of calculating and average.

OpenStudy (anonymous):

I am in seventh grade this is hard

OpenStudy (tkhunny):

You need to know how to calculate an average. Try this one. 4, 6. What is the average of those two numbers? Add them up and divide the sum by the number of elements you added up.

OpenStudy (anonymous):

4+6 =10 10 dived by ?

OpenStudy (♪chibiterasu):

You divide by the total amount of numbers. How many numbers are there?

OpenStudy (tkhunny):

Good start. 4+6 = 10. How many numbers did you add up? 2 10/2 = 5 is the correct answer.

OpenStudy (anonymous):

there is 75,81,88,92

OpenStudy (tkhunny):

Can you calculate the average of those four numbers?

OpenStudy (anonymous):

75+81+88+92 =112 112 / 4 =24

OpenStudy (tkhunny):

I see. There is something VERY SUSPICIOUS with your calculator. That sum is nowhere near 112. Try again. It is MUCH more than that.

OpenStudy (anonymous):

75+81+88+92=336 336/4 =84 ?

OpenStudy (tkhunny):

That's the idea. One thing about an average -- it will NEVER be less than the least value and it will NEVER be greater than the greatest value. When you managed 24 as the average, I knew immediately that it was incorrect, because 24 < 75. This is a good thing to think about. The average is always somewhere between the greatest and least value. 75 < 84 < 92 See how your new average falls in the right place?

OpenStudy (anonymous):

kinda

OpenStudy (tkhunny):

That's a good start. Now, let's think about the problem statement. We just showed that the current average is 84. We need to know what score on the last test will guarantee an average of at least 80. Thinking time.... If the average is already 84, will we have to get more than 84 to keep the average over 80 or can we get less than 84 and still keep the average over 80?

OpenStudy (anonymous):

true

OpenStudy (tkhunny):

What's true? It wasn't a true/false question. Read it again.

OpenStudy (anonymous):

like maybe 80-84 or 84-100

OpenStudy (anonymous):

tkhunny, she is in 7th grade, im in 9th and can hardly undertand the question. Stop making her do what she cant

OpenStudy (tkhunny):

If you score higher than the average, you will make the average increase. Score 85 and the average will go up a little. Score 97 and the average will go up quite a bit. Score 84, and what happens? That is already the average so it doesn't move at all. Score 80 and the average drops a little, but it won't go all the way down to 80. Score 70 and the average will drop more. Will it go below 80? Maybe. Now we know that any score of 80 or above won't hurt. We'll still have an average over 80. How do we find out how far down we can go?

OpenStudy (♪chibiterasu):

Nothing wrong with learning. =)

OpenStudy (tkhunny):

@tijolosoccer Every student will be required to learn how to calculate an average. Did you miss it two years ago? Follow along. See if you learn anything.

OpenStudy (anonymous):

@sam141101

OpenStudy (sam141101):

what?

OpenStudy (tkhunny):

We already know 5+81+88+92=336 336/4 =84 Let's add one more test score and see what happens. 75+81+88+92+78 = 336+78 = 414 414/5 = 82.8 We just did the same process as before. Still following?

OpenStudy (sam141101):

im not good at this stuff

OpenStudy (sam141101):

sorry

OpenStudy (tkhunny):

Then we can continue searching for lower scores until we find the last one that works. 75+81+88+92+76 = 336+76 = 412 412/5 = 82.4 Well, that didn't move much.

OpenStudy (tkhunny):

75+81+88+92+70 = 336+70 = 406 406/5 = 81.2 -- Still over 80, even with a 70 on that last test.

OpenStudy (tkhunny):

75+81+88+92+64 = 336+64 = 400 400/5 = 80.0 -- We found it!! Must get a 64. A 63 will not make it!